OpenStudy (anonymous):

how can we find the domain and range of Sin^{-1} (x-y)

OpenStudy (ranga):

Something is missing in this problem. There should be an equation and there isn't one here. Also is that sine inverse or arc sine as in: \[\Large \sin ^{-1}(x - y)\]?

OpenStudy (anonymous):

yes it is like that

OpenStudy (anonymous):

i wanted to know the method of finding domain and range of inverse functions....

OpenStudy (anonymous):

range is always [-1,1] of arc sin

OpenStudy (anonymous):

but here 2 variables are involved

OpenStudy (anonymous):

not to worry.... arc sin is always confined between -1 and 1 and domain is real

OpenStudy (anonymous):

real no.***

OpenStudy (ranga):

The sine function has a minimum value of -1 and a maximum value of +1. So the (x - y) here has to be confined to [-1, 1]. That is all we can say from the data.

OpenStudy (ranga):

Normally when they ask for domain and range it means what are the allowed values for x and what are the allowed values for y. Here all you can say is: -1 <= (x - y) <= 1

OpenStudy (ranga):

If sin(A) = B then we can be sure B has to be: -1 <= B <= 1 So A = arc sin(B). We can only take arc sin when the value is between [-1, 1] So in this problem: -1 <= (x - y) <= 1 I don't know what else they are expecting.

OpenStudy (anonymous):

thanks...

OpenStudy (ranga):

sure.

OpenStudy (ranga):

 The allowed values for x and y fall in between those two straight lines. So in reality each of x and y can be from -infinity to +infinity as long as -1 <= (x - y) <= 1 So there is a constraint equation. While each of a and y can take any value, as soon as one is defined the other is constrained by -1 <= (x - y) <= 1